# Adapted optimal transport between Gaussian processes in discrete time

### v1

@@ 3. The Knothe-Rosenblatt coupling
This is a special case of Lemma 3.3 below. ∎
Corollary 3.2.
The Cholesky map $A\mapsto L$ is an isometry bewteen the metric spaces $(\mathscr{S}_{++}(N),d_{\mathrm{KR}})$ and $(\mathscr{L}_{++}(N),\|\cdot\|_{\mathrm{F}})$.
The next lemma provides a potential improvement of the Knothe-Rosenblatt coupling and motivates the expression of $\mathcal{AW}_{2}(\mu,\nu)$ in (1.1). It is an elementary version of the control problem in [7, Proposition 2.2] (also see [12]). However, while the synchronous coupling is optimal there, the Knothe-Rosenblatt coupling is not necessarily optimal even when both $\mu$ and $\nu$ are Markovian (and Gaussian).
Lemma 3.3.
Let $\epsilon^{X}=\epsilon^{X}_{1:N}$ and $\epsilon^{Y}=\epsilon^{Y}_{1:N}$ be random vectors such that
(3.5) $\begin{bmatrix}\epsilon^{X}\\ \epsilon^{Y}\end{bmatrix}\sim\mathcal{N}_{2N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}I&P\\ P&I\end{bmatrix}\right),$
where $P_{st}=\delta_{st}\rho_{t}$ ($\delta_{st}$ is the Kronecker delta) and $\rho_{t}\in[-1,1]$ are constants. Let $\pi\in\Pi(\mu,\nu)$ be the distribution of $(X,Y)$ where
[…]
Then $\pi\in\Pi_{bc}(\mu,\nu)$. Among all couplings of this form, the transport cost $\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\pi(\mathrm{d}x,\mathrm{d}y)$ is minimized if and only if
The minimum value is $\|a-b\|^{2}+d_{\mathrm{ABW}}^{2}(A,B)$.
Proof.
By recentering we may assume $a=b=0$. From the criterion below Definition 2.1 we see that $\pi$ is bicausal. Using (3.5), the transport cost under $\pi$ is given by
$\begin{split}&\mathbb{E}_{\pi}\left[\|X-Y\|^{2}\right]=\mathbb{E}_{\pi}\left[\|L\epsilon^{X}-M\epsilon^{Y}\|^{2}\right]\\ &=\mathbb{E}_{\pi}\left[(\epsilon^{X})^{\top}L^{\top}L\epsilon^{X}+(\epsilon^{Y})^{\top}M^{\top}M\epsilon^{Y}-(\epsilon^{X})^{\top}L^{\top}M\epsilon^{Y}-(\epsilon^{Y})^{\top}M^{\top}L\epsilon^{X}\right]\\ &=\operatorname{tr}(A)+\operatorname{tr}(B)-2\sum_{t=1}^{N}\rho_{t}(L^{\top}M)_{tt}.\end{split}$
It is clear that to minimize this quantity we should pick $\rho_{t}$ to match the sign of $(L_{A}^{\top}L_{B})_{tt}$. This yields (3.7) and the desired minimum value. ∎
Note that when $(L_{A}^{\top}L_{B})_{tt}=0$ the transport cost (3.6) is independent of $\rho_{t}$. This suggests that the $\mathcal{AW}_{2}$-optimal coupling is not unique when $\operatorname{diag}(L_{A}^{\top}L_{B})$ has a zero entry. An explicit example is given in Section 6.
@@ 4. Dynamic programming principle
(4.1) $\begin{split}V_{t}(x_{1:t},y_{1:t})=\inf_{\pi_{t+1}}\int_{\mathbb{R}\times\mathbb{R}}V_{t+1}(x_{1:(t+1)},y_{1:(t+1)})\pi_{t+1}(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1}),\end{split}$
@@
The DPP reduces the AOT problem (2.6) to a sequence of one-dimensional OT problems between the conditional marginals. While the DPP is messy for generic distributions, the Gaussian assumption allows us to solve (4.1) explicitly (when the cost is quadratic). In the rest of this section we let $\mu=\mathcal{N}(a,A=LL^{\top})$ and $\nu=\mathcal{N}(b,B=MM^{\top})$ where $L,M\in\mathscr{L}_{++}(N)$. By an adapted Wasserstein coupling $\pi_{\mathrm{AW}}^{\mu,\nu}$ between $\mu$ and $\nu$ we mean a coupling $\pi$ of the form (3.5) where the correlation coefficients satisfy (3.7). This notion is handy when solving the DPP. Observe that since $(L^{\top}M)_{NN}=L_{NN}M_{NN}>0$, in (3.7) we always have $\rho_{NN}=+1$.
@@
(4.2) $(\pi_{\mathrm{AW}}^{\mu,\nu})_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})=\pi_{\mathrm{AW}}^{\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}}),\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})},$
@@
be the block representation of the Cholesky matrix $L$. Then
@@
By Lemma 4.3, we have
$\begin{split}\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}})&=\mathcal{N}_{N-t}(a_{\mathbf{t}^{\prime}}+L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}(x_{\mathbf{t}}-a_{\mathbf{t}}),L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top}),\\ \nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})&=\mathcal{N}_{N-t}(b_{\mathbf{t}^{\prime}}+M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}(x_{\mathbf{t}}-b_{\mathbf{t}}),M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top}).\end{split}$
Under an adapted Wasserstein coupling $\pi_{\mathrm{AW}}^{\mu,\nu}$, write $X=a+L\epsilon^{X}$ and $Y=b+M\epsilon^{Y}$. From (4.5), we have
$\begin{split}X_{{\bf t}^{\prime}}&=a_{\mathbf{t}^{\prime}}+L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}(X_{\mathbf{t}}-a_{\mathbf{t}})+L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}\epsilon_{\mathbf{t}^{\prime}}^{X},\\ Y_{{\bf t}^{\prime}}&=b_{\mathbf{t}^{\prime}}+M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}(Y_{\mathbf{t}}-b_{\mathbf{t}})+M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}\epsilon_{\mathbf{t}^{\prime}}^{Y}.\end{split}$
@@
under the mapping
$(a_{\mathbf{t}^{\prime}}+L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}(x_{\mathbf{t}}-a_{\mathbf{t}})+L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}},b_{\mathbf{t}^{\prime}}+M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}(y_{\mathbf{t}}-b_{\mathbf{t}})+M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}).$
In particular, $(\pi_{\mathrm{AW}}^{\mu,\nu})_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is a coupling of $\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}})$ and $\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})$ and has the form (3.5). A straightforward computation gives
@@
It follows that $\operatorname{diag}(L^{T}M)_{\mathbf{t}^{\prime}}=\operatorname{diag}(L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top}M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}})$. Hence $P_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}$ satisfies (3.7) for the conditional covariance matrices and we conclude that $(\pi_{\mathrm{AW}}^{\mu,\nu})_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is a version of $\pi_{\mathrm{AW}}^{\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}}),\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})}$. ∎
@@
By recentering we may assume $a=b=0$. Thus
$\begin{split}\mathbb{E}_{\mu}[X_{\mathbf{t}^{\prime}}|X_{\mathbf{t}}=x_{\mathbf{t}}]=L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}x_{\mathbf{t}},\quad\mathbb{E}_{\nu}[Y_{\mathbf{t}^{\prime}}|Y_{\mathbf{t}}=y_{\mathbf{t}}]=M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}y_{\mathbf{t}}.\\ \end{split}$
We claim that the value function is given by $V_{N}(x_{1:N},y_{1:N})=\|x_{1:N}-y_{1:N}\|^{2}$ (the terminal value) and, for $0\leq t<N$,
(4.6) $\begin{split}V_{t}(x_{\mathbf{t}},y_{\mathbf{t}})&=\int_{\mathbb{R}^{N-t}\times\mathbb{R}^{N-t}}\|x-y\|^{2}\pi_{\mathrm{AW}}^{\mu,\nu}(\mathrm{d}x_{\mathbf{t}^{\prime}},\mathrm{d}y_{\mathbf{t}^{\prime}}|x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\|x_{\mathbf{t}}-y_{\mathbf{t}}\|^{2}+\|L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}x_{\mathbf{t}}-M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}y_{\mathbf{t}}\|^{2}\\ &\quad+d_{\mathrm{ABW}}^{2}(L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top},M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top}),\end{split}$
where the second inequality follows from Lemma 3.3 and the definition of $\pi_{\mathrm{AW}}^{\mu,\nu}$. [^2] Note that the value of the integral is independent of the version of $\pi_{\mathrm{AW}}^{\mu,\nu}$ used. If so, the conclusion follows by letting $t=0$.
We argue by induction. We use the shorthand $t_{+}=t+1$ and $t_{-}=t-1$ as well as $\mathbf{t}_{+}=1:(t+1)$, $\mathbf{t}_{+}^{\prime}=(t+2):N$ and $\mathbf{t}_{-}^{\prime}=t:N$. Suppose (4.6) holds for $t+1$, so that
@@
Observe that this is a quadratic function in $x_{t+1}$ and $y_{t+1}$. When we integrate it against $\pi_{t+1}(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1})$ for $\pi_{t+1}\in\Pi(\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}}))$, only the integral of the $x_{t+1}y_{t+1}$ term depends on the choice of the coupling.
To obtain the coefficient of $x_{t+1}y_{t+1}$ we begin by analyzing $L_{\mathbf{t}_{++},\mathbf{t}_{+}}^{-1}$ and $M_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}$. Write, in block form,
@@
$\begin{split}L_{\mathbf{t}_{+}^{\prime},\mathbf{t}_{+}}L_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}x_{\mathbf{t}_{+}}&=\begin{bmatrix}L_{\mathbf{t}_{+}^{\prime},\mathbf{t}}&L_{\mathbf{t}_{+}^{\prime},t_{+}}\end{bmatrix}\left[\begin{array}[]{c|c}L_{\mathbf{t},\mathbf{t}}^{-1}&0_{t\times 1}\\ \hline\cr\frac{-1}{L_{t_{+},t_{+}}}L_{t_{+},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}&\frac{1}{L_{t_{+},t_{+}}}\end{array}\right]\begin{bmatrix}x_{\mathbf{t}}\\ x_{t_{+}}\end{bmatrix}\\ &=Cx_{\mathbf{t}}+\frac{1}{L_{t_{+},t_{+}}}L_{\mathbf{t}_{+}^{\prime},t_{+}}x_{t_{+}},\end{split}$
where $C$ is a constant matrix (for $t$ fixed) given by
$C=L_{\mathbf{t}_{+}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}-\frac{1}{L_{t_{+},t_{+}}}L_{\mathbf{t}_{++}^{\prime},t_{+}}L_{t_{+},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}.$
@@
$M_{\mathbf{t}_{+}^{\prime},\mathbf{t}_{+}}M_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}y_{\mathbf{t}_{+}}=Dy_{\mathbf{t}}+\frac{1}{M_{t_{+},t_{+}}}M_{\mathbf{t}_{+}^{\prime},t_{+}}y_{t_{+}},$
where $D$ is defined analogously.
@@
$\begin{split}\alpha_{t+1}&:=1+\frac{1}{L_{t+,t+}M_{t+,t+}}L_{\mathbf{t}_{+}^{\prime},t_{+}}^{\top}M_{\mathbf{t}_{+}^{\prime},t_{+}}=\frac{1}{L_{t+,t+}M_{t+,t+}}(L^{\top}M)_{t+,t+}.\end{split}$
Since $L_{t+,t+}M_{t+,t+}>0$, the sign of $\alpha_{t+1}$ is the same as that of $(L^{\top}M)_{t+,t+}$ or they are both zero. Completing the squares in (4.7), we have
(4.8) $V_{t+1}(x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\ =\|x_{\mathbf{t}}-y_{\mathbf{t}}\|^{2}+(x_{t+1}-\alpha_{t+1}y_{t+1})^{2}\pi(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1})+\cdots,$
@@
(4.9) $\inf_{\pi_{t+1}\in\Pi(\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}}))}\int_{\mathbb{R}\times\mathbb{R}}V_{t+1}(x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\pi(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1}).$
Note that $\mu_{t+1}(\cdot|x_{\mathbf{t}})$ and $\nu_{t+1}(\cdot|y_{\mathbf{t}})$ are univariate Gaussian distributions (and so have finite second moments). From (4.8), $\pi_{t+1}$ is optimal for (4.9) if and only if it is optimal for
$\inf_{\pi\in\Pi(\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}}))}\int_{\mathbb{R}\times\mathbb{R}}(x_{t+1}-\alpha_{t+1}y_{t+1})^{2}\pi(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1}),$

### v4

@@ 3. The Knothe-Rosenblatt coupling
Observe that $\pi^{P}$ is the law of $(X,Y)=(a+L\epsilon^{X},b+M\epsilon^{Y})$, where $\epsilon^{X}=\epsilon_{1:N}^{X}$ and $\epsilon^{Y}=\epsilon_{1:N}^{Y}$ are random vectors with
(3.7) $\begin{bmatrix}\epsilon^{X}\\ \epsilon^{Y}\end{bmatrix}\sim\mathcal{N}_{2N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}I&P\\ P&I\end{bmatrix}\right).$
Since $L,M\in\mathscr{L}_{++}(N)$, from the equivalent condition below Definition 2.1 we see that $\pi$ is a bicausal coupling of $(\mu,\nu)$ (we recover $\pi_{\mathrm{KR}}^{\mu,\nu}$ when $P=I$). Using (3.7), the transport cost under $\pi^{P}$ is given by
$\begin{split}&\mathbb{E}_{\pi^{P}}\left[\|X-Y\|^{2}\right]=\mathbb{E}_{\pi^{P}}\left[\|(a+L\epsilon^{X})-(b+M\epsilon^{Y})\|^{2}\right]\\ &=\|a-b\|^{2}+\mathbb{E}_{\pi^{P}}\left[(\epsilon^{X})^{\top}L^{\top}L\epsilon^{X}+(\epsilon^{Y})^{\top}M^{\top}M\epsilon^{Y}-(\epsilon^{X})^{\top}L^{\top}M\epsilon^{Y}-(\epsilon^{Y})^{\top}M^{\top}L\epsilon^{X}\right]\\ &=\|a-b\|^{2}+\operatorname{tr}(A)+\operatorname{tr}(B)-2\sum_{t=1}^{N}\rho_{t}(L^{\top}M)_{tt}.\end{split}$
It is clear that to minimize this quantity we should pick $\rho_{t}$ to match the sign of $(L^{\top}M)_{tt}$. This yields (3.6) and the desired minimum value. ∎
In Lemma 3.3 the correlations $\rho_{t}$ are deterministic. In principle, we may consider more general bicausal couplings where $\rho_{t}$ is previsible, i.e., $\rho_{t}=\rho_{t}(X_{1:(t-1)},Y_{1:(t-1)})$. It turns out that this is not necessary in our discrete time Gaussian setting. We also observe that when $(L^{\top}M)_{tt}=0$ the transport cost $\mathbb{E}_{\pi^{P}}\left[\|X-Y\|^{2}\right]$ does not depend on $\rho_{t}$. This suggests that the $\mathcal{AW}_{2}$-optimal coupling is not unique when $\operatorname{diag}(L^{\top}M)$ contains a zero entry. An explicit example is given in Section 6.
@@ 4. Dynamic programming principle
(4.1) $\begin{split}V_{t}(x_{1:t},y_{1:t})=\inf_{\pi_{t+1}}\int_{\mathbb{R}\times\mathbb{R}}V_{t+1}(x_{1:(t+1)},y_{1:(t+1)})\pi_{t+1}(\mathrm{d}x_{t+1}\mathrm{d}y_{t+1}),\end{split}$
@@
The DPP reduces the AOT problem (2.6) to a sequence of one-dimensional OT problems between the conditional marginals. While the DPP is complicated for generic distributions, the Gaussian assumption allows us to solve (4.1) explicitly when the cost is quadratic. In the rest of this section we fix $\mu=\mathcal{N}(a,A=LL^{\top})$ and $\nu=\mathcal{N}(b,B=MM^{\top})$ where $L,M\in\mathscr{L}_{++}(N)$. By an adapted Wasserstein coupling $\pi_{\mathrm{AW}}^{\mu,\nu}$ of $(\mu,\nu)$, we mean a coupling $\pi^{P}$ in the context of Lemma 3.3 where $P\in\mathscr{P}(N)$ satisfies (3.6); it is unique if and only if $(L^{\top}M)_{tt}\neq 0$ for all $t\in[N]$. This notion is handy when solving the DPP. Observe that since $(L^{\top}M)_{NN}=L_{NN}M_{NN}>0$, in (3.6) we always have $\rho_{NN}=+1$.
@@
$(\pi_{\mathrm{AW}}^{\mu,\nu})_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})=\pi_{\mathrm{AW}}^{\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}}),\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})},$
@@
be the block diagonal representation of the Cholesky matrix $L$. Then
@@
Under an adapted Wasserstein coupling $\pi_{\mathrm{AW}}^{\mu,\nu}$, write $X=a+L\epsilon^{X}$ and $Y=b+M\epsilon^{Y}$ where $(\epsilon^{X},\epsilon^{Y})$ is distributed as (3.7). From (4.4), we have
$\begin{split}X_{{\bf t}^{\prime}}=a_{\mathbf{t}^{\prime}}+L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}(X_{\mathbf{t}}-a_{\mathbf{t}})+L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}\epsilon_{\mathbf{t}^{\prime}}^{X}\quad\text{and}\quad Y_{{\bf t}^{\prime}}=b_{\mathbf{t}^{\prime}}+M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}(Y_{\mathbf{t}}-b_{\mathbf{t}})+M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}\epsilon_{\mathbf{t}^{\prime}}^{Y}.\end{split}$
@@
under the affine map
$(\epsilon_{\mathbf{t}^{\prime}}^{X},\epsilon_{\mathbf{t}^{\prime}}^{Y})\mapsto(a_{\mathbf{t}^{\prime}}+L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}(x_{\mathbf{t}}-a_{\mathbf{t}})+L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}\epsilon_{\mathbf{t}^{\prime}}^{X},b_{\mathbf{t}^{\prime}}+M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}(y_{\mathbf{t}}-b_{\mathbf{t}})+M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}\epsilon_{\mathbf{t}^{\prime}}^{Y}).$
In particular, $(\pi_{\mathrm{AW}}^{\mu,\nu})_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is a coupling of $(\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}}),\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}}))$ and has the form (3.5). A straightforward computation gives
@@
It follows that $\operatorname{diag}(L^{T}M)_{\mathbf{t}^{\prime}}=\operatorname{diag}(L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top}M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}})$. Hence $P_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}$ satisfies (3.6) for the conditional covariance matrices and we conclude that $(\pi_{\mathrm{AW}}^{\mu,\nu})_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is an adapted Wasserstein coupling between $\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}})$ and $\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})$. ∎
@@
By recentering (i.e., considering $X-a$ and $Y-a$) we may assume $a=b=0$. From Lemma 4.3, we have
(4.5) $\begin{split}\mathbb{E}_{\mu}[X_{\mathbf{t}^{\prime}}|X_{\mathbf{t}}=x_{\mathbf{t}}]=L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}x_{\mathbf{t}}\quad\text{and}\quad\mathbb{E}_{\nu}[Y_{\mathbf{t}^{\prime}}|Y_{\mathbf{t}}=y_{\mathbf{t}}]=M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}y_{\mathbf{t}}.\\ \end{split}$
Let $V_{t}(x_{1:t},y_{1:t})$ be the value function in Theorem 4.1. We claim that
$V_{N}(x_{1:N},y_{1:N})=\|x_{1:N}-y_{1:N}\|^{2}$
(the terminal value) and, for $0\leq t<N$,
(4.6) $\begin{split}&V_{t}(x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int_{\mathbb{R}^{N-t}\times\mathbb{R}^{N-t}}\|x-y\|^{2}\pi_{\mathrm{AW}}^{\mu,\nu}(\mathrm{d}x_{\mathbf{t}^{\prime}}\mathrm{d}y_{\mathbf{t}^{\prime}}|x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\|x_{\mathbf{t}}-y_{\mathbf{t}}\|^{2}+\|L_{\mathbf{t}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}x_{\mathbf{t}}-M_{\mathbf{t}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}y_{\mathbf{t}}\|^{2}+d_{\mathrm{ABW}}^{2}(L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}L_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top},M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}M_{\mathbf{t}^{\prime},\mathbf{t}^{\prime}}^{\top}),\end{split}$
where the second equality follows from Lemma 3.3 and the definition of $\pi_{\mathrm{AW}}^{\mu,\nu}$. [^4] From Lemma 3.3, the value of the integral is independent of the version of $\pi_{\mathrm{AW}}^{\mu,\nu}$ used. If so, the conclusion follows by letting $t=0$.
We argue by induction. We use the shorthand $t_{+}=t+1$ as well as $\mathbf{t}_{+}=1:(t+1)$ and $\mathbf{t}_{+}^{\prime}=(t+2):N$. Suppose (4.6) holds for $t+1$, so that
@@
Observe that this is a quadratic function of $x_{t+1}$ and $y_{t+1}$ which appear in the second and third terms. When we integrate it against $\pi_{t+1}\in\Pi(\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}}))$, only the integral of the $x_{t+1}y_{t+1}$ term depends on the choice of the coupling.
To obtain the coefficient of $x_{t+1}y_{t+1}$ we begin by analyzing $L_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}$ and $M_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}$. Write, in block form,
@@
(4.8) $\begin{split}L_{\mathbf{t}_{+}^{\prime},\mathbf{t}_{+}}L_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}x_{\mathbf{t}_{+}}&=\begin{bmatrix}L_{\mathbf{t}_{+}^{\prime},\mathbf{t}}&L_{\mathbf{t}_{+}^{\prime},t_{+}}\end{bmatrix}\left[\begin{array}[]{c|c}L_{\mathbf{t},\mathbf{t}}^{-1}&0_{t\times 1}\\ \hline\cr\frac{-1}{L_{t_{+},t_{+}}}L_{t_{+},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}&\frac{1}{L_{t_{+},t_{+}}}\end{array}\right]\begin{bmatrix}x_{\mathbf{t}}\\ x_{t_{+}}\end{bmatrix}\\ &=\left(L_{\mathbf{t}_{+}^{\prime},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}-\frac{1}{L_{t_{+},t_{+}}}L_{\mathbf{t}_{+}^{\prime},t_{+}}L_{t_{+},\mathbf{t}}L_{\mathbf{t},\mathbf{t}}^{-1}\right)x_{\mathbf{t}}+\frac{1}{L_{t_{+},t_{+}}}L_{\mathbf{t}_{+}^{\prime},t_{+}}x_{t_{+}}.\end{split}$
@@
$M_{\mathbf{t}_{+}^{\prime},\mathbf{t}_{+}}M_{\mathbf{t}_{+},\mathbf{t}_{+}}^{-1}y_{\mathbf{t}_{+}}=\left(M_{\mathbf{t}_{+}^{\prime},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}-\frac{1}{M_{t_{+},t_{+}}}M_{\mathbf{t}_{+}^{\prime},t_{+}}M_{t_{+},\mathbf{t}}M_{\mathbf{t},\mathbf{t}}^{-1}\right)y_{\mathbf{t}}+\frac{1}{M_{t_{+},t_{+}}}M_{\mathbf{t}_{+}^{\prime},t_{+}}y_{t_{+}}.$
@@
$\begin{split}\alpha_{t+1}&:=1+\frac{1}{L_{t_{+},t_{+}}M_{t_{+},t_{+}}}L_{\mathbf{t}_{+}^{\prime},t_{+}}^{\top}M_{\mathbf{t}_{+}^{\prime},t_{+}}=\frac{1}{L_{t_{+},t_{+}}M_{t_{+},t_{+}}}(L^{\top}M)_{t_{+},t_{+}}.\end{split}$
Since $L_{t_{+},t_{+}}M_{t_{+},t_{+}}>0$, the sign of $\alpha_{t+1}$ is the same as that of $(L^{\top}M)_{t+1,t+1}$ or they are both zero. Completing the squares in (4.7), we have
(4.9) $V_{t+1}(x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\ =\|x_{\mathbf{t}}-y_{\mathbf{t}}\|^{2}+(x_{t+1}-\alpha_{t+1}y_{t+1})^{2}+\cdots,$
@@
(4.10) $\inf_{\pi_{t+1}\in\Pi(\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}}))}\int_{\mathbb{R}\times\mathbb{R}}V_{t+1}(x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\pi(\mathrm{d}x_{t+1}\mathrm{d}y_{t+1}),$
where $\mu_{t+1}(\cdot|x_{\mathbf{t}})$ and $\nu_{t+1}(\cdot|y_{\mathbf{t}})$ are univariate Gaussian distributions. From (4.9), $\pi_{t+1}$ is optimal for (4.10) if and only if it is optimal for
$\inf_{\pi\in\Pi(\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}}))}\int_{\mathbb{R}\times\mathbb{R}}(x_{t+1}-\alpha_{t+1}y_{t+1})^{2}\pi(\mathrm{d}x_{t+1}\mathrm{d}y_{t+1}),$